# Time-Series Foundation Model for Value-at-Risk

### v2

@@
Abstract
This study is the first to explore the application of a time-series foundation model for Value-at-Risk (VaR) estimation. Foundation models, pre-trained on vast and varied datasets, can be used in a zero-shot setting with relatively minimal data or further improved through finetuning. We compare the performance of Google’s model, called TimesFM, against conventional parametric and non-parametric models, including GARCH, Generalized Autoregressive Score (GAS), and empirical quantile estimates, using daily returns from the S&P 100 index and its constituents over 19 years. Our backtesting results indicate that in terms of the actual-over-expected ratio the fine-tuned TimesFM model consistently outperforms traditional methods. Regarding the quantile score loss function, it achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the best or among the top performers in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. We also found that fine-tuning significantly improves the results, and the model should not be used in zero-shot settings.
keywords:
Value at Risk , Foundation Models , TimesFM , Time series forecasting
journal: SSRN
@@ Abstract
[inst1]organization=Tampere University, Department of Computing Sciences, Financial Computing and Data Analytics Group,addressline=Kalevantie 4, city=Tampere, postcode=33100, country=Finland
@@
[inst2]organization=IIT Ropar, Department of Mathematics,addressline=Ropar, city=Punjab, postcode=140001, country=India
@@ 1 Introduction
While the application of Machine Learning (ML) in various fields in financial market research has gained significant attention—such as in asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging (Gu et al., 2020; Christensen et al., 2023; Tran et al., 2018; Mikkilä and Kanniainen, 2023)—the risk management literature still largely relies on traditional econometric techniques, especially in the area of Value-at-Risk (VaR). One reason for this may be that Deep Learning (DL) models require substantial amounts of data, which poses a significant limitation when applying them to calculate risk metrics for individual stocks. VaR estimations are often based on daily return observations (for example, see González-Rivera et al., 2004; Patton et al., 2019; Taylor, 2019, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, machine learning has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are built on standard deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al., 2021). In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al., 2024), Time-GPT (Garza and Mergenthaler-Canseco, 2024), and LagLlama (Rasul et al., 2023).
@@
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data (Tran et al., 2019; Guijarro-Ordonez et al., 2021; Chen et al., 2024). Historically, methods like GARCH-based models have been effective for capturing dependencies and volatility clustering in market data Hoga and Demetrescu (2023). Recent approaches, however, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), extend these capabilities by generating realistic synthetic data for distribution forecasting (Nareklishvili et al., 2023; Athey et al., 2021; Hofert et al., 2022). These generative models allow for flexible, non-parametric modeling of data structures that were once challenging to approximate with traditional econometric tools. Ericson et al. (2024) provide a comprehensive comparative review, highlighting that deep generative methods outperform traditional benchmarks in capturing complex financial dynamics for VaR estimation. These advancements underscore the potential of AI-driven methods in risk forecasting. Our approach diverges by utilizing a time-series foundation model, pre-trained on extensive and various data sets, to predict VaR. Compared to traditional supervised ML models, which are trained for specific tasks, these time-series forecasting models are pre-trained on vast, diverse datasets containing billions of real-world time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The word ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM, for VaR estimation and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. Additionally, this paper seeks to determine how much fine-tuning enhances accuracy in the context of VaR predictions. This paper could employ alternative foundation models as several have been published recently, and more are on the way. However, we decided to focus on the use TimesFM because it is an open-source [^1] https://github.com/google-research/TimesFM., modifiable, and free-of-charge, [^2] As an additional note, the authors of Google’s TimesFM empirically demonstrated that their model outperformed TimeGPT in forecasting accuracy. making our results completely replicable. [^3] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. Moreover, our goal is not to find the best possible foundation model among those available, as they are likely to be modified and entirely new ones can be published on a monthly basis, but rather to address more fundamental question if the state-of-the-art parametric and semi-parametric approaches can be outperformed.
@@
We conduct an extensive empirical evaluation of the performance of the foundation time-series model with daily return data on the 91 constituents of the S&P 100 index over 19 years. For each constituent, we calculate 1-day VaR on a daily basis. The foundation model is applied both with and without fine-tuning to understand how much the results improve with fine-tuning, which has certain computational cost. The fine-tuning is performed on a daily basis using a rolling window technique.
@@
Our main finding is that, in terms of the actual-over-expected ratio, the fine-tuned time-series foundation model outperforms existing econometric approaches for VaR estimation. Specifically, the fine-tuned foundation model consistently demonstrates superior performance, on average, across all 91 constituents of the S&P 100 included in the analysis. Regarding the quantile score loss function, which was the second backtesting metric, the fine-tuned foundation model achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the top performer or among the best in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Moreover, we found that fine-tuning is practically necessary for the foundation model; using it as pre-trained in zero-shot settings results in clearly worse performance compared to the best econometric approaches. We consider this as a remarkable result on the performance of Google’s TimesFM model as this model is not explicitly designed for this purpose and was not even pre-trained with stock return data. This problem is statistically challenging due to its focus on heteroscedastic processes with time-varying conditional volatility. In this light, the paper is valuable not only for financial risk management but also as a benchmark to test AI time-series models against state-of-the-art statistical approaches on a highly complex issue.
@@ 2.1 Time-Series Foundation Models
TimesFM model (Das et al., 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focuse solely on output generation processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco, 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason the encoder component does not add significant value.
@@
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition (Makridakis et al., 2020; Spiliotis et al., 2020). However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy
@@ 2.2 Benchmark Models for VaR
  • 2.
    ARMA-GARCH Models (Angelidis et al., 2004):
@@
    3. (c)
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle, 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • 3.
    One factor Gas model (Patton et al., 2019): This model belongs to the class of CAViaR models (Taylor, 2005), where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^4] GAS model was originally proposed in Creal et al. (2013). Patton et al. (2019) proposed a dynamic GAS model employing the class of loss functions proposed in Fissler and Ziegel (2016), called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in Patton et al. (2019). The evolution of VaR for the Gas 1 factor model is given
@@ 3 Data and Experiments
We now apply the models discussed in the last section to the forecasting of VaR for daily returns on the S&P 100 index along with its 91 constituents. The data used in our analysis was obtained from Thomson Reuters Datastream, covering a period of approximately 19 years, from January 2005 to September 30, 2023, resulting in 4,876 observations. Throughout this period, data on every day was available for 91 stocks. [^5] The S&P100 Global Index typically comprises 100 companies; however, as of October 1, 2023, it included 102 stocks (as Roche Holding AG and GOOGLE have two classes of shares). However, due to the unavailability of data for certain stocks, specifically Synlogic Inc (SYBX UR Equity), Alphabet Inc (GOOG UW Equity), Seven & i Holdings Co Ltd (3382 JT Equity), Texas Instruments Inc (TXN UW Equity), Shell PLC (SHEL LN Equity), Philip Morris International Inc (PM UN Equity), Mastercard Inc (MA UN Equity), Broadcom Inc (AVGO UW Equity), Engie SA (ENGI FP Equity), PepsiCo Inc (PEP UW Equity), and Honeywell International Inc (HON UW Equity)—the analysis was conducted on the remaining 91 constituents. The daily return for $i$-th stock at $t$-th day is calculated as $r_{it}=\frac{p_{it}-p_{it-1}}{p_{it-1}};$ $i=1,\ldots,n,\ t=1,\ldots,T,$ where $p_{it}$ and $p_{it-1}$ are respectively, the closing prices at $t$-th and $(t-1)$-th day. We divide the data into two consecutive non-overlapping samples while maintaining the temporal ordering of the data. We use the first ten years of data, i.e., January 2005 to December 2014, for estimating model parameters and the remaining 9 years, i.e., January 2015 to September 2023, for testing the performance of various models.
@@ 3.1 Experiment Setup
We implement the recently introduced pre-trained time-series model, Google’s TimesFM, for forecasting Value at risk and compare its performance against state-of-the-art parametric and non-parametric approaches, namely, GARCH model with the standardized residual modelled as empirical distribution function (G-EDF), normal distribution (G-N) and Student’s-t (G-t) distribution, one factor Generalized AutoRegressive Score model (GAS) and rolling window method (Historical).
@@
TimesFM is originally trained on a large number of public and synthetic datasets and exhibits robust out-of-the-box zero-shot performance in comparison to the accuracy of various state-of-the-art forecasting models specific to individual datasets under consideration. The model is flexible enough to forecast with different input length, prediction length and time granularities at inference time. The pre-trained model returns the point forecasts along with deciles, i.e., the quantiles [^6] As per Google’e TimesFM github page, “We experimentally offer quantile heads but they have not been calibrated after pretraining.” at $10\%,20\%,\ldots,90\%$ levels. We experimented with different prediction horizons with a fixed input length of 2 years (512 days). Specifically, we set the prediction horizon to 1 day, 21 days and 63 days and denote these models as PT1, PT21 and PT63. For PT1, the input window shifts daily with the prediction points, while for PT21 and PT63, we use a fixed set of input observations for daily forecasts over the subsequent 21 and 63 days, respectively. For example, with PT21 (or PT63), we estimated one-day VaRs over a 21-day (or 63-day) period starting from January 2, 2015, using the most recent 512 observations. After this, we shifted the training data forward by 21 days (or 63 days) to incorporate more recent data and applied the pre-trained model for the next 21 (or 63) days, generating one-day VaRs estimates over this period. The variation in prediction horizon allows us to examine the model’s ability to generalize across forecasting horizons of varying lengths, i.e., short and long horizons. However, as mentioned above, the pre-trained models, PT1, PT21 and PT63, only produced VaR forecast at 90% level.
@@
In the fine-tuning process, we divided the first ten-year period, from January 2005 to December 2014, into two subsets: the training set from January 2005 to December 2011, and the validation set from January 2012 to December 2014. The validation data was used for hyperparameter optimization and to track the model’s performance after every epoch. That is, validation data is used to tune hyper-parameters by evaluating the performance on validation set during training, and selecting the best model check-point, that is, the one with lowest validation loss. With a fixed input length of 512, we fine-tuned the pre-trained model separately for three different prediction horizons. Specifically, we obtained the fine-tuned model FT1 using a prediction horizon of 1 day, FT21 with a horizon of 21 dats, and FT63 with a horizon of 63 days. Overall, three fine-tuned models were generated, denoted as FT1, FT21, and FT63, corresponding to the different prediction horizons. With fine-tuning we re-trained the model for the quantile levels 0.01, 0.025, 0.05, and 0.1 as the pre-trained models used quantiles 0.1, 0.2, … 0.9 only.

### v3

@@
Time-Series Foundation Model for Value-at-Risk
[…]
Time-Series Foundation Model for Value-at-Risk
[…]
This study is the first to explore the performance of a time-series foundation model for Value-at-Risk (VaR) estimation. Foundation models, pre-trained on vast and varied datasets, can be used in a zero-shot setting with relatively minimal data or further improved through finetuning. We compare the performance of Google’s model, called TimesFM, against conventional parametric and non-parametric models, including GARCH, Generalized Autoregressive Score (GAS), and empirical quantile estimates, using daily returns from the S&P 100 index and its constituents over 19 years. Our backtesting results indicate that in terms of the actual-over-expected ratio, the fine-tuned TimesFM model consistently outperforms traditional methods. Regarding the quantile score loss function, it achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the best or among the top performers in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Fine-tuning significantly improves accuracy, indicating that zero-shot use is not optimal for VaR estimation.
[…]
Value at Risk, AI, Foundation models, TimesFM, Time series forecasting
@@ 1 Introduction
While Machine Learning (ML) has significantly advanced financial market research—such as in asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging (Gu et al. 2020, Christensen et al. 2023, Tran et al. 2018, Mikkilä and Kanniainen 2023)—the risk management literature still largely relies on traditional econometric techniques, especially in the area of Value-at-Risk (VaR). One reason for this may be that Deep Learning (DL) models require substantial amounts of data, which poses a significant limitation when applying them to calculate risk metrics for individual stocks. VaR estimations are often based on daily return observations (for example, see González-Rivera et al. 2004, Patton et al. 2019, Taylor 2019, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, ML has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are based on deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al. 2021). In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al. 2024), Time-GPT (Garza and Mergenthaler-Canseco 2024), and LagLlama (Rasul et al. 2023).
@@
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data (Tran et al. 2019, Guijarro-Ordonez et al. 2021, Chen et al. 2024). Historically, methods like GARCH-based models have been effective for capturing dependencies and volatility clustering in market data (Hoga and Demetrescu 2023). Recent approaches, however, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), enhance distribution forecasting by generating realistic synthetic data (Nareklishvili et al. 2023, Athey et al. 2021, Hofert et al. 2022). Such models provide flexible, non-parametric approaches for data structures that traditional econometric tools may struggle to approximate. Ericson et al. (2024) presents a comparative review showing that deep generative methods outperform traditional benchmarks in modeling complex financial dynamics for VaR estimation. These advancements highlight the potential of AI-driven approaches in risk forecasting. This study stands out from previous research by adopting a time-series foundation model, pre-trained on diverse datasets, to forecast VaR. Unlike traditional supervised ML models trained for specific tasks, time-series foundation models are pre-trained on vast datasets spanning billions of time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The term ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM, for VaR estimation and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. An additional objective is to assess the extent to which fine-tuning improves the accuracy of VaR predictions. While several foundation models have been recently introduced, this study focuses on TimesFM due to its open-source availability [^1] https://github.com/google-research/TimesFM., modifiability, and free accessibility, ensuring the results are fully replicable. [^2] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. Furthermore, the authors of TimesFM have empirically demonstrated that their model outperforms TimeGPT in forecasting accuracy, adding further motivation for its selection. Rather than identifying the best foundation model, which is subject to frequent updates and new developments, this study aims to address the fundamental question of whether state-of-the-art parametric and semi-parametric methods can be outperformed.
@@
We conduct an extensive empirical evaluation of the performance of the foundation time-series model using daily return data on the 91 constituents of the S&P 100 index over 19 years. For each constituent, we calculate 1-day VaR on a daily basis. The foundation model is applied both with and without fine-tuning to understand how much the results improve with fine-tuning, which has certain computational cost. The fine-tuning is performed on a daily basis using a rolling window technique.
@@
Our main finding is that, in terms of the actual-over-expected ratio, the fine-tuned time-series foundation model outperforms existing econometric approaches for VaR estimation. Specifically, the fine-tuned foundation model consistently demonstrates superior performance, on average, across all 91 constituents of the S&P 100 included in the analysis. Regarding the quantile score loss function, which is the second backtesting metric, the fine-tuned foundation model achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the top performer or among the best in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Moreover, we found that fine-tuning is practically necessary for the foundation model; using it as pre-trained in zero-shot settings results in significantly worse performance compared to the best econometric approaches. We consider this as a remarkable result on the performance of Google’s TimesFM model as this model is not explicitly designed for this purpose and was not even pre-trained with stock return data. This problem is statistically challenging due to its focus on heteroscedastic processes with time-varying conditional volatility. In this light, the paper is valuable not only for financial risk management, but also as a benchmark for testing AI time-series models against state-of-the-art statistical approaches on a highly complex issue.
@@ 2.1 Time-Series Foundation Models
TimesFM model (Das et al. 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focus solely on output generation processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason, the encoder component does not add significant value.
@@
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition (Makridakis et al. 2020, Spiliotis et al. 2020). However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy.
@@ 2.2 Benchmark Models for VaR
  • ARMA-GARCH Models (Angelidis et al. 2004):
@@
    3. 3.
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • One factor Gas model (Patton et al. 2019): This model belongs to the class of CAViaR models (Taylor 2005), where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^3] GAS model was originally proposed in Creal et al. (2013). Patton et al. (2019) proposed a dynamic GAS model employing the class of loss functions proposed in Fissler and Ziegel (2016), called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in Patton et al. (2019). The evolution of VaR for the Gas 1 factor model is given
@@ 3 Data and Experiments
We now apply the models discussed in the last section to the forecasting of VaR for daily returns on the S&P 100 index along with its 91 constituents. The data used in our analysis was obtained from Thomson Reuters Datastream, covering a period of approximately 19 years, from January 2005 to September 30, 2023, resulting in 4,876 observations. Throughout this period, data on every day was available for 91 stocks. [^4] The S&P100 Global Index typically comprises 100 companies; however, as of October 1, 2023, it included 102 stocks (as Roche Holding AG and GOOGLE have two classes of shares). However, due to the unavailability of data for certain stocks, specifically Synlogic Inc (SYBX UR Equity), Alphabet Inc (GOOG UW Equity), Seven & i Holdings Co Ltd (3382 JT Equity), Texas Instruments Inc (TXN UW Equity), Shell PLC (SHEL LN Equity), Philip Morris International Inc (PM UN Equity), Mastercard Inc (MA UN Equity), Broadcom Inc (AVGO UW Equity), Engie SA (ENGI FP Equity), PepsiCo Inc (PEP UW Equity), and Honeywell International Inc (HON UW Equity)—the analysis was conducted on the remaining 91 constituents. The daily return for $i$-th stock at $t$-th day is calculated as $r_{it}=\frac{p_{it}-p_{it-1}}{p_{it-1}};$ $i=1,\ldots,n,\ t=1,\ldots,T,$ where $p_{it}$ and $p_{it-1}$ are respectively, the closing prices at $t$-th and $(t-1)$-th day. We divide the data into two consecutive, non-overlapping samples while maintaining the temporal ordering of the data. We use the first ten years of data, i.e., January 2005 to December 2014, for estimating model parameters and the remaining 9 years, i.e., January 2015 to September 2023, for testing the performance of various models.
@@ 3.1 Experiment Setup
We implement the recently introduced pre-trained time-series model, Google’s TimesFM, for forecasting Value at risk and compare its performance against state-of-the-art parametric and non-parametric approaches, namely, GARCH model with the standardized residual modeled as empirical distribution function (G-EDF), normal distribution (G-N) and Student’s-t (G-t) distribution, one factor Generalized AutoRegressive Score model (GAS) and rolling window method (Historical).
@@
TimesFM is originally trained on a large number of public and synthetic datasets and exhibits robust out-of-the-box zero-shot performance in comparison to the accuracy of various state-of-the-art forecasting models specific to individual datasets under consideration. The model is flexible enough to forecast with different input length, prediction length and time granularities at inference time. The pre-trained model returns the point forecasts along with deciles, i.e., the quantiles [^5] As per Google’e TimesFM github page, “We experimentally offer quantile heads but they have not been calibrated after pretraining.” at $10\%,20\%,\ldots,90\%$ levels. We experimented with different prediction horizons with a fixed input length of 2 years (512 days). Specifically, we set the prediction horizon to 1 day, 21 days and 63 days and denote these models as PT1, PT21, and PT63. For PT1, the input window shifts daily with the prediction points, while for PT21 and PT63, we use a fixed set of input observations for daily forecasts over the subsequent 21 and 63 days, respectively. For example, with PT21 (or PT63), we estimated one-day VaRs over a 21-day (or 63-day) period starting from January 2, 2015, using the most recent 512 observations. After this, we shifted the training data forward by 21 days (or 63 days) to incorporate more recent data and applied the pre-trained model for the next 21 (or 63) days, generating one-day VaRs estimates over this period. The variation in prediction horizon allows us to examine the model’s ability to generalize across forecasting horizons of varying lengths, i.e., short and long horizons. However, as mentioned above, the pre-trained models, PT1, PT21, and PT63, only produced VaR forecast at 90% level.
@@
In the fine-tuning process, we divided the first ten-year period, from January 2005 to December 2014, into two subsets: the training set from January 2005 to December 2011, and the validation set from January 2012 to December 2014. The validation data was used for hyperparameter optimization and to track the model’s performance after every epoch. That is, validation data is used to tune hyper-parameters by evaluating the performance on validation set during training, and selecting the best model check-point, that is, the one with the lowest validation loss. With a fixed input length of 512, we fine-tuned the pre-trained model separately for three different prediction horizons. Specifically, we obtained the fine-tuned model FT1 using a prediction horizon of 1 day, FT21 with a horizon of 21 days, and FT63 with a horizon of 63 days. Overall, three fine-tuned models were generated, denoted as FT1, FT21, and FT63, corresponding to the different prediction horizons. With fine-tuning, we re-trained the model for the quantile levels 0.01, 0.025, 0.05, and 0.1 as the pre-trained models used quantiles 0.1, 0.2,… 0.9 only.